### v2

@@ 3.1 Regularity conditions and their relationship
    $f(x)+\langle g,y-x\rangle+\frac{\mu_{\mathrm{s}}}{2}\|y-x\|^{2}$ $\leq f(y),$ (SC)
    $\forall x,y\in[f\leq f^{\star}+\nu],g\in\hat{\partial}f(x).$
@@
    $\mu_{\mathrm{r}}\cdot\mathrm{dist}^{2}(x,S)$ $\leq\mathopen{}\mathclose{{}\left\langle g,x-\hat{x}}\right\rangle,$ (RSI)
    $\forall x\in[f\leq f^{\star}+\nu],g\in\hat{\partial}f(x),\hat{x}\in\Pi_{S}(x).$
@@
    Error bound (EB) [^1] Error bound is closely related to the notion of metric subregularity at an optimal point $x^{\star}$ for $0\in\hat{\partial}f(\mathbb{R}^{n})$ [17, Definition 2.3]: there exist a constant $\alpha>0$ and a set $\mathcal{U}$ containing $x^{\star}$ such that $\mathrm{dist}(x,(\hat{\partial}f)^{-1}(0))\leq\alpha\cdot\mathrm{dist}(0,\hat{\partial}f(x)),\forall x\in\mathcal{U}$.: there exists a constant $\mu_{\mathrm{e}}>0$ such that
@@
    Polyak-Łojasiewicz (PL) inequality [^2] To be consistent with the smooth case in [10], we call the property PL Polyak-Łojasiewicz, which is usually used for smooth functions. The property PL is actually a special case of the Kurdyka-Łojasiewicz inequality $\varphi^{\prime}(f(x)-f^{\star})\mathrm{dist}(0,\hat{\partial}f(x))\geq 1$ with $\varphi(s)=cs^{1/2}$ and $c>0$.: there exists a constant $\mu_{\mathrm{p}}>0$ such that
@@
All the regularity conditions above are defined over a sublevel set $[f\leq f^{\star}+\nu]$. If $v=+\infty$, then they are global. In particular, SC imposes a quadratic lower bound for every point in the sublevel set. On the other hand, RSI, EB and PL all require a certain growth of the minimal norm element in ^$\hat{\partial}f(x)$ when moving away from its solution set $S$ or optimal value $f^{\star}$. It is easy to see that RSI, EB and PL all imply that every stationary point $0\in\hat{\partial}f(x)$ in the sublevel set $[f\leq f^{\star}+\nu]$ is a global minimum (but they do not imply the uniqueness of stationary points). Finally, QG shows that $f(x)$ grows at least quadratically when moving away from the solution set $S$. Unlike RSI, EB and PL, this QG alone allows suboptimal stationary points.
@@
RSI $\mu_{\mathrm{r}}$ $[\frac{1}{\mu_{\mathrm{r}}},+\infty)$ $(0,\frac{\mu_{\mathrm{r}}^{2}}{2\mu_{\mathrm{r}}+\rho}]$ $(0,\frac{\mu_{\mathrm{r}}}{2}]$
@@
PL $(0,\frac{\mu_{\mathrm{p}}-2\rho}{4}]$ $[\frac{1}{\mu_{\mathrm{p}}},+\infty)$ $\mu_{\mathrm{p}}$ $(0,\frac{\mu_{\mathrm{p}}}{2}]$
QG $(0,\frac{\mu_{\mathrm{q}}-\rho}{2}]$ $[\frac{2}{\mu_{\mathrm{q}}-\rho},+\infty)$ $(0,\frac{(\mu_{\mathrm{q}}-\rho)^{2}}{4\mu_{\mathrm{q}}}]$ $\mu_{\mathrm{q}}$
Note that if a function $f$ is convex and satisfies QG with $\mu_{\mathrm{q}}>0$, then 8 holds naturally as $f$ is naturally 00-weakly convex. Theorem 3.1 includes [10, Theorem 2] as a special case, in which only $L$-smooth functions are considered. It is easy to see that all $L$-smooth functions are also $L$-weakly convex. Even for differentiable functions, Theorem 3.1 is more general than [10, Theorem 2] in the sense that 1) we require no Lipschitz constant $L$ for gradients to ensure the equivalency among EB, PL, and QG in the convex case; 2) the condition $\mu_{\mathrm{q}}>\rho$ is new and does not mean that $f$ is convex. We will provide further illustrative examples in Section 3.2. Under the condition $\mu_{\mathrm{q}}>\rho$, we can further infer the coefficients for different regularity conditions from each other; the details are listed in Table 1. The proof of Theorem 3.1 relies heavily on the notion of slope in definition 2.1, Ekeland’s variational principle in Theorem 2.2, and the technical result in Lemma 2.2. We will present the proof details in Section 3.3. Alternative proofs based on subgradient flows [18] are also possible. Indeed, one key step in the proof of [10, Theorem 2] is based on gradient flows for smooth functions, which is a special case of subgradient flows for nonsmooth cases.
Remark 3.1.
The regularity conditions EB, QG and PL have been discussed for different function classes in the literature. For the smooth case, we refer to [10, Theorem 2] for a nice summary; also see [23, 22] for related discussions. For nonsmooth convex functions, the equivalence between EB and QG has been recognized in [8, Theorem 3.3] and [17, Theorem 3.3], and the equivalence between PL and QG is established in [18, Theorem 5]. Thus, EB, PL, and QG are equivalent for the class of nonsmooth convex functions [19, Proposition 2]; see also [20] for a recent discussion. Our Theorem 3.1 extends these results to $\rho$-weakly convex functions. The most closely related work is [21] which focuses on nonsmooth optimization using Taylor-like models. Indeed, we specialize the proof of [21, Theorem 3.7, Proposition 3.8, and Corollary 5.7] into our setting and prove the relationship in 8 and 7 directly using the slope technique. We note that the implication from QG to EB/PL is not true in general. Yet, with the condition $\mu_{\mathrm{q}}>\rho$ in Theorem 3.1, all four regularity conditions RSI, EB, PL and QG are equivalent. $\square$
@@ 3.2 Examples
In principle, all the five properties in Theorem 3.1 are generalizations of quadratic functions to non-quadratic, nonconvex, and even nonsmooth cases. For illustration, let us first consider the simplest quadratic function $f(x)=x^{2}$, which is convex and differentiable. It is clear that ^$\hat{\partial}f(x)=\{2x\}$ and $S=\{0\}$. It is also immediate to verify that SC holds with $0<\mu_{\mathrm{s}}\leq 2$, RSI holds with $0<\mu_{\mathrm{r}}\leq 2$, EB holds with $\mu_{\mathrm{e}}\geq 1/2$, PL holds with $0<\mu_{\mathrm{p}}\leq 2$, and QG holds with $0<\mu_{\mathrm{q}}\leq 2$. Consider another simple convex function $f(x)=x^{2},\text{if}\leavevmode\nobreak\ |x|\leq 1,$ and $f(x)=\frac{1}{2}x^{4}+\frac{1}{2}$ otherwise. All the five properties hold for this function, but it is not $L$-smooth globally.
@@
$f(x)=\begin{cases}-x^{2}+1&\text{if}\leavevmode\nobreak\ -1<x<-0.5,\\ 3(x+1)^{2}&\text{otherwise}.\end{cases}$
It can be verified that the function is not convex (due to the part $-1<x<-0.5$) but $2$-weakly convex with a unique minimizer $x^{\star}=-1$ and the minimum value $f^{\star}=0$. See Figure 1 for an illustration of these properties. The Fréchet subdifferential of this function can be calculated as
^$\hat{\partial}f(x)=\begin{cases}-2x&\text{if}\leavevmode\nobreak\ -1<x<-0.5,\\ 6(x+1)&\text{if}\leavevmode\nobreak\ x<-1\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ x>-0.5,\\ [0,2]&\text{if}\leavevmode\nobreak\ x=-1,\\ [1,3]&\text{if}\leavevmode\nobreak\ x=-0.5.\end{cases}$
@@
$\mathrm{dist}(0,\hat{\partial}f(x))=\begin{cases}|2x|&\text{if}\leavevmode\nobreak\ -1<x<-0.5,\\ |6(x+1)|&\text{if}\leavevmode\nobreak\ x<-1\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ x>-0.5,\\ 0&\text{if}\leavevmode\nobreak\ x=-1,\\ 1&\text{if}\leavevmode\nobreak\ x=-0.5.\end{cases}$
As expected from Theorem 3.1, we can see that QG holds for $0<\mu_{\mathrm{q}}\leq 6$, EB holds for $\mu_{\mathrm{e}}\geq 1/2$, and PL holds for $30<\mu_{\mathrm{p}}\leq 2/$ 3. These can also be observed in Figure 1.
@@
The directions SC $\Rightarrow$ RSI, RSI $\Rightarrow$ EB, EB $\Rightarrow$ PL, and QG with $\mu_{\mathrm{q}}>\rho$Rightarrow RSI follow easily from some elementary algebras. The directions PL $\Rightarrow$ EB and EB $\Rightarrow$ PL require the equivalence in Lemma 2.1. Most directions are not difficult to prove, except that the direction PL $\Rightarrow$ EB requires some sophisticated arguments involving Lemmas 2.2 and 2.2. To start the proof, we first recall that $S$ denotes the set of minimizers of $f$ and $f^{\star}=\min_{x\in\mathbb{R}^{n}}f(x)$ and we assume $S$ is nonempty.
@@
    $f(x)+\mathopen{}\mathclose{{}\left\langle g,\hat{x}-x}\right\rangle+\frac{\mu_{\mathrm{s}}}{2}\cdot\mathrm{dist}^{2}(x,S)\leq f(\hat{x})$
@@
  • RSI $\Rightarrow$ EB: Suppose $f$ satisfies RSI with constant $\mu_{\mathrm{r}}>0$. Let $x\in[f\leq f^{\star}+\nu]$, $g$ be the minimal norm element in ^$\hat{\partial}f(x)$, and $\hat{x}\in\Pi_{S}(x)$. Then, by definition of RSI, we have
@@
  • EB $\Rightarrow$ PL: Suppose $f$ satisfies EB with constant $\mu_{\mathrm{e}}>0$. Let $x\in[f\leq f^{\star}+\nu]$, $\hat{x}\in\Pi_{S}(x)$, and $g$ be the minimal norm element in ^$\hat{\partial}f(x)$. From the subdifferential property for $\rho$-weakly convex function (i.e. 26 in Appendix A), we have
@@
  • PL $\Rightarrow$ EB: This direction requires the introduction of limiting subdifferential of a function $h:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ at $\bar{x}$ defined as [37, Equation 18]
    $\partial h(\bar{x})=\{s\in\mathbb{R}^{n}\mid\exists x_{k}\to\bar{x}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ s_{k}\in\hat{\partial}h(x_{k})\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ s_{k}\to s\}.$
    Note that we follow the standard convention, using $\partial h$ to denote both the convex subdifferential and limiting subdifferential of $h$. Suppose $f$ satisfies PL with constant $\mu_{\mathrm{p}}>0$. As $f$ is $\rho$-weakly convex, applying Lemma 2.1 on PL shows that
@@
    Let $g(x)=(f(x)-f^{\star})^{1/2}$. For all $\bar{x}\in[0<g\leq\sqrt{\nu}]$, we have the following inequalities
    $|\nabla g|(\bar{x})$ $\geq\liminf_{x\to\bar{x},g(x)\to g(\bar{x})}|\nabla g|(x)\overset{(b)}{=}\mathrm{dist}(0,\partial g(\bar{x}))$
    $\overset{(c)}{\geq}\mathrm{dist}(0,\frac{1}{2}(f(\bar{x})-f^{\star})^{-1/2}\partial f(\bar{x}))$
    =$\overset{(d)}{=}\mathrm{dist}(0,\frac{1}{2}(f(\bar{x})-f^{\star})^{-1/2}\hat{\partial}f(\bar{x}))$
    $=\frac{1}{2}(f(\bar{x})-f^{\star})^{-1/2}\mathrm{dist}(0,\hat{\partial}f(\bar{x}))$
    =$\overset{(e)}{=}\frac{1}{2}(f(\bar{x})-f^{\star})^{-1/2}|\nabla f|(\bar{x})$
    $\overset{(f)}{\geq}\frac{1}{2}(f(\bar{x})-f^{\star})^{-1/2}\cdot\sqrt{2\mu_{\mathrm{p}}}(f(\bar{x})-f^{\star})^{1/2}$
    $=\sqrt{\frac{\mu_{\mathrm{p}}}{2}},$
[…]
    Fix a point $x\in[0<g\leq\sqrt{\nu}]$. We choose $K>g(x)\sqrt{\frac{2}{\mu_{\mathrm{p}}}}$. It is obvious that we have $g(x)<K\sqrt{\frac{\mu_{\mathrm{p}}}{2}}$ and
    $|\nabla g|(u)$ $\geq\sqrt{\frac{\mu_{\mathrm{p}}}{2}},\quad\forall u\in[0<g\leq g(x),\|x-u\|\leq K.$
@@
    $\leq\frac{1}{\mu_{\mathrm{p}}}|\nabla f|(x)=\frac{1}{\mu_{\mathrm{p}}}\mathrm{dist}(0,\hat{\partial}f(x)),\forall x\in[f\leq f^{\star}+\nu]$
    where the first inequality is from Lemma 2.2, the second inequality comes from 9, and the third inequality applies Lemma 2.1. This completes the proof of PL $\Rightarrow$ EB.
@@
  • QG with $\mu_{\mathrm{q}}>\rho$Rightarrow RSI: Let $x\in[f\leq f^{\star}+\nu]$, $\hat{x}\in\Pi_{S}(x)$, and $g\in\hat{\partial}f(x)$. From the assumption of QG and the property of subgradient of weakly convex function in 26, we have

### v3

@@ 3.1 Regularity conditions and their relationship
    $f(x)+\langle g,y-x\rangle+\frac{\mu_{\mathrm{s}}}{2}\|y-x\|^{2}$ $\leq f(y),\quad\forall x,y\in[f\leq f^{\star}+\nu],g\in\hat{\partial}f(x).$ (SC)
@@
    $\mu_{\mathrm{r}}\cdot\mathrm{dist}^{2}(x,S)$ $\leq\mathopen{}\mathclose{{\left\langle g,x-\hat{x}}}\right\rangle,\quad\forall x\in[f\leq f^{\star}+\nu],g\in\hat{\partial}f(x),\hat{x}\in\Pi_{S}(x).$ (RSI)
@@
    Error bound (EB) [^2] Error bound is closely related to the notion of metric subregularity at an optimal point $x^{\star}$ for $0\in\hat{\partial}f(x^{\star})$ [artacho2008characterization, Definition 2.3]: there exist a constant $\alpha>0$ and a set $\mathcal{U}$ containing $x^{\star}$ such that $\mathrm{dist}(x,(\hat{\partial}f)^{-1}(0))\leq\alpha\cdot\mathrm{dist}(0,\hat{\partial}f(x)),\forall x\in\mathcal{U}$.: there exists a constant $\mu_{\mathrm{e}}>0$ such that
@@
    Polyak-Łojasiewicz (PL) inequality [^3] To be consistent with the smooth case in [karimi2016linear], we call the property PL Polyak-Łojasiewicz, which is usually used for smooth functions. The property PL is actually a special case of the Kurdyka-Łojasiewicz inequality $\varphi^{\prime}(f(x)-f^{\star})\mathrm{dist}(0,\hat{\partial}f(x))\geq 1$ with $\varphi(s)=cs^{1/2}$ and $c>0$.: there exists a constant $\mu_{\mathrm{p}}>0$ such that
@@
All the regularity conditions above are defined over a sublevel set $f\leq f^{\star}+\nu]$. If $v=+\infty$, then they are global. In particular, SC imposes a quadratic lower bound at every point in the sublevel set. On the other hand, RSI, EB and PL all require a certain growth of the minimal norm element in $\hat{\partial}f(x)$ when moving away from its solution set $S$ or optimal value $f^{\star}$. It is easy to see that RSI, EB and PL all imply that every stationary point $0\in\hat{\partial}f(x)$ in the sublevel set $f\leq f^{\star}+\nu]$ is a global minimum (but they do not imply the uniqueness of stationary points). Finally, QG shows that $f(x)$ grows at least quadratically when moving away from the solution set $S$. Unlike RSI, EB and PL, this QG alone allows suboptimal stationary points.
@@
RSI $\mu_{\mathrm{r}}$ $\frac{1}{\mu_{\mathrm{r}}},+\infty)$ $(0,\frac{\mu_{\mathrm{r}}^{2}}{2\mu_{\mathrm{r}}+\rho}]$ $(0,\frac{\mu_{\mathrm{r}}}{2}]$
@@
PL $(0,\frac{\mu_{\mathrm{p}}-2\rho}{4}]$ $\frac{1}{\mu_{\mathrm{p}}},+\infty)$ $\mu_{\mathrm{p}}$ $(0,\frac{\mu_{\mathrm{p}}}{2}]$
QG $(0,\frac{\mu_{\mathrm{q}}-\rho}{2}]$ $\frac{2}{\mu_{\mathrm{q}}-\rho},+\infty)$ $(0,\frac{(\mu_{\mathrm{q}}-\rho)^{2}}{4\mu_{\mathrm{q}}}]$ $\mu_{\mathrm{q}}$
Note that if a function $f$ is convex and satisfies QG with $\mu_{\mathrm{q}}>0$, then 8 holds naturally as $f$ is naturally 0-weakly convex. Theorem 3.1 includes [karimi2016linear, Theorem 2] as a special case, in which only $L$-smooth functions are considered. It is easy to see that all $L$-smooth functions are also $L$-weakly convex. Even for differentiable functions, Theorem 3.1 is more general than [karimi2016linear, Theorem 2] in the sense that 1) we require no Lipschitz constant $L$ for gradients to ensure the equivalency among EB, PL, and QG in the convex case; 2) the condition $\mu_{\mathrm{q}}>\rho$ is new and does not mean that $f$ is convex. We will provide further illustrative examples in Section 3.2. Under the condition $\mu_{\mathrm{q}}>\rho$, we can further infer the coefficients for different regularity conditions from each other; the details are listed in Table 1.
The proof of Theorem 3.1 relies heavily on the notion of slope in Definition 1, Ekeland’s variational principle in Theorem 2.2, and the technical result in Lemma 2.2. We will present the proof details in Section 3.3. Alternative proofs based on subgradient flows [bolte2017error] are also possible. Indeed, one key step in the proof of [karimi2016linear, Theorem 2] is based on gradient flows for smooth functions, which is a special case of subgradient flows for nonsmooth cases.
Remark 1.
The regularity conditions EB, QG and PL have been discussed for different function classes in the literature. For the smooth case, we refer to [karimi2016linear, Theorem 2] for a nice summary; also see [rebjock2023fast, guille2021study] for related discussions. For nonsmooth convex functions, the equivalence between EB and QG has been recognized in [drusvyatskiy2018error, Theorem 3.3] and [artacho2008characterization, Theorem 3.3], and the equivalence between PL and QG is established in [bolte2017error, Theorem 5]. Thus, EB, PL, and QG are equivalent for the class of nonsmooth convex functions [ye2021variational, Proposition 2]; see also [zhu2023unified] for a recent discussion. Our Theorem 3.1 extends these results to $\rho$-weakly convex functions. The most closely related work is [drusvyatskiy2021nonsmooth] which focuses on nonsmooth optimization using Taylor-like models. Indeed, we specialize the proof of [drusvyatskiy2021nonsmooth, Theorem 3.7, Proposition 3.8, and Corollary 5.7] into our setting and prove the relationship in 8 and 7 directly using the slope technique. We note that the implication from QG to EB/PL is not true in general. Yet, with the condition $\mu_{\mathrm{q}}>\rho$ in Theorem 3.1, all four regularity conditions RSI, EB, PL and QG are equivalent. $\square$
@@ 3.2 Examples
In principle, all the five properties in Theorem 3.1 are generalizations of quadratic functions to non-quadratic, nonconvex, and even nonsmooth cases. For illustration, let us first consider the simplest quadratic function $f(x)=x^{2}$, which is convex and differentiable. It is clear that $\hat{\partial}f(x)=\{2x\}$ and $S=\{0\}$. It is also immediate to verify that SC holds with $0<\mu_{\mathrm{s}}\leq 2$, RSI holds with $0<\mu_{\mathrm{r}}\leq 2$, EB holds with $\mu_{\mathrm{e}}\geq 1/2$, PL holds with $0<\mu_{\mathrm{p}}\leq 2$, and QG holds with $0<\mu_{\mathrm{q}}\leq 2$. Consider another simple convex function $f(x)=x^{2},\text{if}\penalty 10000\ |x|\leq 1,$ and $f(x)=\frac{1}{2}x^{4}+\frac{1}{2}$ otherwise. All the five properties hold for this function, but it is not $L$-smooth globally.
@@
[…]
It can be verified that the function is not convex (due to the part 1<x<0.5-1<x<-0.5) but $2$-weakly convex with a unique minimizer $x^{\star}=-1$ and the minimum value $f^{\star}=0$. See Figure 1 for an illustration of these properties. The Fréchet subdifferential of this function can be calculated as
[…]
@@
[…]
As expected from Theorem 3.1, we can see that QG holds for $0<\mu_{\mathrm{q}}\leq 6$, EB holds for $\mu_{\mathrm{e}}\geq 1/2$, and PL holds for $0<\mu_{\mathrm{p}}\leq 2/3$. These can also be observed in Figure 1.
@@
The directions SC $\Rightarrow$ RSI, RSI $\Rightarrow$ EB, EB $\Rightarrow$ PL, and QG with $\mu_{\mathrm{q}}>\rho$ $\Rightarrow$ RSI follow easily from some elementary algebras. The directions PL $\Rightarrow$ EB and EB $\Rightarrow$ PL require the equivalence in Lemma 2.1. Most directions are not difficult to prove, except that the direction PL $\Rightarrow$ EB requires some sophisticated arguments involving Lemmas 2.2 and 2.2. To start the proof, we first recall that $S$ denotes the set of minimizers of $f$ and $f^{\star}=\min_{x\in\mathbb{R}^{n}}f(x)$ and we assume $S$ is nonempty.
@@
[…]
@@
  • RSI $\Rightarrow$ EB: Suppose $f$ satisfies RSI with constant $\mu_{\mathrm{r}}>0$. Let $x\in[f\leq f^{\star}+\nu]$, $g$ be the minimal norm element in $\hat{\partial}f(x)$, and $\hat{x}\in\Pi_{S}(x)$. Then, by definition of RSI, we have
@@
  • EB $\Rightarrow$ PL: Suppose $f$ satisfies EB with constant $\mu_{\mathrm{e}}>0$. Let $x\in[f\leq f^{\star}+\nu]$, $\hat{x}\in\Pi_{S}(x)$, and $g$ be the minimal norm element in $\hat{\partial}f(x)$. From the subdifferential property for $\rho$-weakly convex function (i.e. 29 in Appendix A), we have
@@
  • PL $\Rightarrow$ EB: Suppose $f$ satisfies PL with constant $\mu_{\mathrm{p}}>0$. As $f$ is $\rho$-weakly convex, applying Lemma 2.1 on PL shows that
@@
    Define the function $g$ as $x\mapsto g(x)=(f(x)-f^{\star})^{1/2}$. We claim that
    $|\nabla g|(x)\geq\sqrt{\frac{\mu_{\mathrm{p}}}{2}},\;\forall x\in[0<g\leq\sqrt{\nu}].$ (10)
    With 10, we can establish the implication PL $\Rightarrow$ EB as follows. Fix a point $x\in[0<g\leq\sqrt{\nu}]$ and choose $K>g(x)\sqrt{\frac{2}{\mu_{\mathrm{p}}}}$. We have $g(x)<K\sqrt{\frac{\mu_{\mathrm{p}}}{2}}$ and
    $|\nabla g|(u)$ $\geq\sqrt{\frac{\mu_{\mathrm{p}}}{2}},\quad\forall u\in[0<g\leq g(x)],\|x-u\|\leq K.$
@@
    $\leq\frac{1}{\mu_{\mathrm{p}}}|\nabla f|(x)=\frac{1}{\mu_{\mathrm{p}}}\mathrm{dist}(0,\hat{\partial}f(x)),\forall x\in[f\leq f^{\star}+\nu],$
    where the first inequality is from Lemma 2.2, the second inequality comes from 9, and the third equality applies Lemma 2.1. This shows the implication PL $\Rightarrow$ EB.
    Proof of the claim 10: Fix any $\bar{x}\in[0<g\leq\sqrt{\nu}]$. We shall show
    $|\nabla g|(\bar{x})\geq\sqrt{\frac{\mu_{\mathrm{p}}}{2}}.$ (11)
    Note that if $|\nabla g|(\bar{x})=+\infty$, then (11) holds trivially. Hence, we only prove (11) for the case $|\nabla g|(\bar{x})<+\infty$. By the definition of the slope $|\nabla f|(\bar{x})=\lim\sup_{y\rightarrow\bar{x}}\frac{(f(\bar{x})-f(y))^{+}}{\lVert y-\bar{x}\rVert}$ and by our choice of $\bar{x}$ satisfying $|\nabla f|(\bar{x})\geq 2\mu_{p}(f(\bar{x})-f^{\star})>0$, we claim that there exists a sequence $\{y_{i}\}$ satisfying the following conditions:
    $\{y_{i}\}\rightarrow\bar{x},\quad\{f(y_{i})\}\uparrow f(\bar{x}),\quad|\nabla f|(\bar{x})=\lim_{i\to+\infty}\frac{f(\bar{x})-f(y_{i})}{\lVert y_{i}-\bar{x}\rVert},$ (12)
    where $\{f(y_{i})\}\uparrow f(\bar{x})$ means $f(y_{i})\leq f(\bar{x})$ for all $i$ and $\{f(y_{i})\}\rightarrow f(\bar{x})$. If 12 holds, then (11) can be readily established using basic calculus. Indeed, using the Taylor expansion of the function $x\mapsto\sqrt{x}$ near $f(\bar{x})-f^{\star}$, we have
    $|\nabla g|(\bar{x})$ $=\limsup_{y\to\bar{x}}\frac{\mathopen{}\mathclose{{\left((f(\bar{x})-f^{\star})^{\frac{1}{2}}-(f(y)-f^{*})^{\frac{1}{2}}}}\right)^{+}}{\|y-\bar{x}\|}$
    $\geq\limsup_{i\rightarrow+\infty}\frac{(f(\bar{x})-f^{*})^{\frac{1}{2}}-(f(y_{i})-f^{*})^{\frac{1}{2}}}{\lVert y_{i}-\bar{x}\rVert}$
    $=\limsup_{i\rightarrow+\infty}\frac{\frac{1}{2(f(\bar{x})-f^{*})^{\frac{1}{2}}}(f(\bar{x})-f(y_{i}))+o(|f(y_{i})-f(\bar{x})|)}{\lVert y_{i}-\bar{x}\rVert}$
    $\overset{(a)}{=}\frac{|\nabla f|(\bar{x})}{2(f(\bar{x})-f^{*})^{\frac{1}{2}}}$
    $\overset{(b)}{\geq}\sqrt{\frac{\mu_{\mathrm{p}}}{2}},$
    where $(a)$ is due to (12), and $(b)$ applies (9). It remains to prove the claim 12.
    Proof of the claim 12: The conditions $\{y_{i}\}\rightarrow\bar{x}$, $f(y_{i})\leq f(\bar{x})$ for all $i$, and $|\nabla f|(\bar{x})=\lim_{i\to+\infty}\frac{f(\bar{x})-f(y_{i})}{\lVert y_{i}-\bar{x}\rVert}$ follow from the definition of $|\nabla f|(\bar{x})$ and the assumption $|\nabla f|(\bar{x})>0$. Below, we further establish $\{f(y_{i})\}\uparrow f(\bar{x})$. Suppose for contradiction that the sequence $\{f(y_{i})\}$ does not converge to $f(\bar{x})$. Then, let $\epsilon=\frac{1}{2}(f(\bar{x})-f^{\star})>0$, there exists a subsequence $\{f(y_{i_{k}})\}_{k}$ such that $\epsilon\leq f(\bar{x})-f(y_{i_{k}})$. However, this subsequence will result in
    $|\nabla g(\bar{x})|$ $\geq\limsup_{k\rightarrow+\infty}\frac{(f(\bar{x})-f^{*})^{\frac{1}{2}}-(f(y_{i_{k}})-f^{*})^{\frac{1}{2}}}{\lVert y_{i_{k}}-\bar{x}\rVert}$
    $\geq\limsup_{k\rightarrow+\infty}\frac{(f(\bar{x})-f^{*})^{\frac{1}{2}}-(f(\bar{x})-f^{*}-\epsilon)^{\frac{1}{2}}}{\lVert y_{i_{k}}-\bar{x}\rVert}$
    $=\limsup_{k\rightarrow+\infty}\frac{(f(\bar{x})-f^{*})^{\frac{1}{2}}-(\frac{1}{2}(f(\bar{x})-f^{*}))^{\frac{1}{2}}}{\lVert y_{i_{k}}-\bar{x}\rVert}$
    $=+\infty,$
    violating the assumption $|\nabla g|(\bar{x})<+\infty$. This proves the claim 12.
@@
  • QG with $\mu_{\mathrm{q}}>\rho$ $\Rightarrow$ RSI: Let $x\in[f\leq f^{\star}+\nu]$, $\hat{x}\in\Pi_{S}(x)$, and $g\in\hat{\partial}f(x)$. From the assumption of QG and the property of subgradient of weakly convex function in 29, we have